OpenStudy (anonymous):

18 5/12 + 8 10/17 + 23 7/12 + 28 7/17 A. 77 B. 77 1/2 C. 78 1/2 D. 79

OpenStudy (plasmataco):

...

OpenStudy (plasmataco):

79.

OpenStudy (plasmataco):

8 10/17 + 28 7/17=37. 18 5/12 + 23 7/12=42. 37+42=79.